# The Negation Of Singer’s Conjecture For The Sixth Algebraic Transfer

### v3

@@ Abstract
Let $\mathscr{A}$ be the Steenrod algebra over the field of characteristic two, $\mathbb{F}_{2}.$ Denote by $GL(q)$ the general linear group of rank $q$ over $\mathbb{F}_{2}.$ The algebraic transfer, introduced by W. Singer [Math. Z. 202 (1989), 493-523], is a rather effective tool for unraveling the intricate structure of the (mod-2) cohomology of the Steenrod algebra, ${\rm Ext}_{\mathscr{A}}^{q,*}(\mathbb{F}_{2},\mathbb{F}_{2}).$ The Kameko homomorphism is one of the useful tools to study the dimension of the domain of the Singer transfer. Singer conjectured that the algebraic transfer is always a monomorphism, but this remains open for all homology degrees $q\geq 5.$ In this paper, by constructing a novel algorithm implemented in the computer algebra system OSCAR for computing $GL(q)$-invariants of the kernel of the Kameko homomorphism, we disprove Singer’s conjecture for bidegree $(6,6+36).$
@@
Adams spectral sequences; Steenrod algebra; Hit problem; Algebraic transfer.
MSC:
[2010] 55Q45, 55S10, 55S05, 55T15.
@@ 1 Introduction and statement of the main outcome
Introduction. Write $\mathbb{F}_{2}$ as the prime field of 2 elements, we use $H^{*}$ (resp. $H_{*}$) to denote $H^{*}(-;\mathbb{F}_{2})$ (resp. $H_{*}(-;\mathbb{F}_{2})$). It is well-known that the Steenrod algebra $\mathscr{A}$ over $\mathbb{F}_{2}$ is the algebra of all stable cohomology operations and plays a vital role in Algebraic Topology. We want to emphasize its importance in stable homotopy theory. One of the most outstanding and essential problems is the computation of stable homotopy groups of spheres. Although many profound results have been obtained, it is a fundamental and challenging problem, and we still need to complete it. Many great mathematicians developed deep theories and practical tools to compute and understand them. One of them is the Adams spectral sequence. It approximates the 2-primary stable homotopy groups of the sphere spectrum $\mathbb{S}^{0}$ starting with the cohomology of the Steenrod algebra, ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}):=\bigoplus_{r\geq 0}{\rm Ext}^{q,r}_{\mathscr{A}}(\widetilde{H}^{*}(\mathbb{S}^{0}),\mathbb{F}_{2})=\bigoplus_{r\geq 0}{\rm Ext}^{q,r}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}).$ Here $q$ is the homological degree and $r$ is associated with the degree in $\mathscr{A}.$ Thus, the computation of ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is significant in stable homotopy theory. For further understanding of Ext groups and related topics, readers may refer, for instance, to the papers authored by Palmieri [20], Wen-Hsiung Lin [14], Chen [4, 5], and Weinan Lin [15, 16].
Despite their considerable import, Ext groups, especially ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ for the Adams spectral sequence, proffer calculative challenges across sundry degrees. In recent times, momentous headway has been accomplished via modern computer algebra systems using minimal resolutions (see also [15, 16] and references therein). Within the scope of this paper, another efficient instrument that we are especially interested in is the Singer algebraic transfer proposed by Singer himself in 1989 [32]. In particular, in the surprising work [32], Singer employs the modular invariant theory to investigate Ext groups through his algebraic transfer homomorphism. Before delving into the details of the Singer transfer and ensuring its accessibility for readers, we would like to recall some pertinent aspects.
@@
Let $\mathcal{V}^{q}$ denote a $q$-dimensional $\mathbb{F}_{2}$-vector space. Since $\mathbb{F}_{2}$ is a prime field of size two, $\mathcal{V}^{q}$ can be regarded as a rank-$q$ elementary abelian 2-group. It is well-known that $H^{*}(\mathcal{V}^{q})\cong S(\mathcal{V}_{*}^{q})$, the symmetric algebra of the dual space $\mathcal{V}_{*}^{q}\equiv H^{1}(\mathcal{V}^{q})$. We can choose $x_{1},x_{2},\ldots,x_{q}$ to be a basis of $H^{1}(\mathcal{V}^{q})$. In this case, $P_{q}:=H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}[x_{1},x_{2},\ldots,x_{q}]$, the connected $\mathbb{N}$-graded polynomial algebra on generators of degree $1,$ equipped with the canonical unstable algebra structure over $\mathscr{A}.$ By dualizing, the mod-2 homology $H_{*}(\mathcal{V}^{q})$ is a divided power algebra on $q$ generators. Let $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q})$ be the subspace of $H_{*}(\mathcal{V}^{q})$ consisting of all elements that are annihilated by all positive degree Steenrod operations. The group $GL(q)$ acts regularly on $\mathcal{V}^{q}$ and therefore on $P_{q}$ and $H_{*}(\mathcal{V}^{q}).$ This action commutes with that of the algebra $\mathscr{A}$ and so acts on $\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}$ and $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q}).$ With the notion that the structure of the Ext groups can be studied through modular invariant theory, Singer [32] formulated a homomorphism denoted as:
@@
There exists a profound association between the domain of $Tr_{q}(\mathbb{F}_{2})$ and the structure of the tensor product $\mathbb{F}_{2}\otimes_{\mathscr{A}}H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}.$ Indeed, we give $\mathbb{F}_{2}$ the trivial $\mathscr{A}$-module structure. That is, the unit in $\mathscr{A}$ acts as a unit, while $Sq^{k}(\mathbb{F}_{2})=0$ for any $k\geq 1.$ Let $\mathscr{A}^{>0}$ denote the positive degree part of $\mathscr{A},$ and put
@@
where $\mathscr{A}^{>0}\cdot P_{q}$ refers to the subspace of $P_{q}$ that comprises all homogeneous polynomials in the form $\sum_{k\geq 1}Sq^{k}(f_{k}),$ with $Sq^{k}\in\mathscr{A}^{>0}$ and $f_{k}\in P_{q}.$ Note that
@@
To provide readers with the context related to the inequality above, we restate the following important fact, which has been mentioned in [31, Remark 2.3]: A key element in analyzing $(QP_{q})_{n}$ is the set of ”spike” monomials—monomials where every exponent is of the form $2^{m}-1$. A classical result establishes that spike monomials do not belong to $\mathscr{A}^{>0}\cdot P_{q}$ (see also [37]). Additionally, in [18], Mothebe constructed a rather involved recursive function to enumerate all spike monomials of degree $n=2^{q-1}-q$, namely
@@
Grasping the structure and executing computations of the dimensions of $(QP_{q})_{n}$ and the invariant $[(QP_{q})_{n}]^{GL(q)}=[(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})))_{n}]^{*}$ stand as exceptionally difficult undertakings, if not impossibilities, even with the help of computer algebra instruments. An additional insightful detail concerning the graded vector space $QP_{q}$ is provided by the Peterson conjecture [21], which was subsequently proven correct by Wood [39]. This conjecture states that $QP_{q}$ is trivial in degrees $n$ if $\mu(n)>q,$ where $\mu(n)$ denotes the minimal integer $\zeta$ for which $n$ can be written as $\sum_{1\leq j\leq\zeta}(2^{d_{j}}-1)$ for some positive integers $d_{j}.$ After considering the aforementioned events, our focus can now shift towards exploring the domain of the algebraic transfer, $\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}H_{n}(\mathcal{V}^{q})$, under the condition $\mu(n)\leq q.$ It is significant to mention that the constraint $\mu(n)\leq q$ has an equivalent useful formulation as $\alpha(n+q)\leq q$, with $\alpha(k)$ being the number of 1’s in the binary expression of the natural number $k$. This facilitates characterizing the pertinent ”families” of $n$ satisfying $\mu(n)\leq q.$
Over the past nearly four decades, the Singer transfer and related aspects have been extensively studied by numerous authors (see, e.g., [1, 3, 6, 7, 8, 9, 10, 11, 12, 13, 17, 19, 21, 22, 33, 34, 35, 23, 24, 25, 26, 27, 28, 29, 30, 31, 36, 37, 38, 39]). In particular, in [1], Boardman showed that $Tr_{3}(\mathbb{F}_{2})$ is also an isomorphism. Remarkably, Singer [32] proved that the algebraic transfer fails to be surjective in bidegree $(5,14),$ after which he proposed the following ensuing unsettled conjecture.
@@
To prove Theorem 1.4, we construct and implement a new algorithm in the OSCAR computer algebra system [40]. The algorithm computes an explicit basis for both the kernel of the Kameko homomorphism and the space $(QP_{q})_{n},$ as well as their corresponding invariants, for any $q$ and $n$ where $n-q$ is even. (Our previous algorithm in [30], implemented in SageMath, did not perform these basis and invariant computations for the kernel of the Kameko homomorphism.) Our reasoning for choosing OSCAR over SageMath for this implementation is detailed in Note 3.4(C) of Section 3. Furthermore, we used this new algorithm to verify previously known results, including those we computed by hand and those published by other authors (see, e.g., [3, 10, 26, 34, 36]). Our algorithm’s output is consistent with these established findings.
@@
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [35], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand and, unfortunately, its correctness has not yet been verified using a computer algebra system (e.g., OSCAR, Magma, or SageMath). In fact, verifying Sum’s counterexample [35] on a computer algebra system is also very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [29]). The difficulty in verification is precisely why Conjecture 1.1 had effectively remained open for all homology degrees $q\geq 5$ until we now refute it by Theorems 1.2, 1.4 and Corollary 1.3 for the case $(q,n)=(6,36),$ and the result has been fully verified using computer algebra systems.
@@
Adopting an alternative perspective, Hưng [10] proposed an interesting notion concerning a critical element that exists within ${\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$. Specifically, a non-zero element $\mathbf{u}$ in ${\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is deemed critical if it satisfies two conditions: firstly, $\mu(2n+q)=q$, and secondly, the image of $\mathbf{u}$ under the classical squaring operation $Sq^{0}$ is zero.
It is well-established that $Sq^{0}$ is a monomorphism in positive stems of ${\rm Ext}_{\mathscr{A}}^{q,q+n}(\mathbb{F}_{2},\mathbb{F}_{2})$ for $q<5,$ thereby implying the absence of any critical element for $q<5.$ Remarkably, Hưng’s work [10, Theorem 5.9] states that Singer’s Conjecture 1.1 is not valid, if the algebraic transfer detects the critical elements.
@@
According to Hưng [12], the algebraic Kahn-Priddy homomorphism
@@
is an epimorphism from ${\rm Im}(Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}))$ onto ${\rm Im}(Tr_{q}(\mathbb{F}_{2}))={\rm Im}(Tr_{q}(\widetilde{H}^{*}\mathbb{S}^{0}))$ in positive stem $n.$ Hence, for $q=5,\,n=26,$ there exists the non-zero element $\widehat{D_{2}}\in{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ with $t_{*}(\widehat{D_{2}})=D_{2}\in{\rm Ext}^{6,6+26}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}),$ which is the image of the transfer
$Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty}):(\mathbb{F}_{2}\otimes_{GL(5)}\mathcal{P}_{\mathscr{A}}(H_{*}\mathcal{V}^{5}\otimes\widetilde{H}_{*}\mathbb{R}P^{\infty}))_{26}\longrightarrow{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2}).$
Consequently, by a completely analogous argument to that used for the elements $\widehat{Ph_{1}}$ and $\widehat{Ph_{2}}$ in [12], it may be concluded that $\widehat{D_{2}}$ is also a critical element and is not in the image of $Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$.
@@
Using Theorem 1.4 and our algorithm given in [30] for determining preimages in the lambda algebra, we hope that there will be an answer to Conjecture 1.7.

### v6

@@ Abstract
Let $\mathscr{A}$ be the Steenrod algebra over the field of characteristic two, $\mathbb{F}_{2}.$ Denote by $GL(q)$ the general linear group of rank $q$ over $\mathbb{F}_{2}.$ The algebraic transfer, introduced by W. Singer [Math. Z. 202 (1989), 493-523], is a rather effective tool for unraveling the intricate structure of the (mod-2) cohomology of the Steenrod algebra, ${\rm Ext}_{\mathscr{A}}^{q,*}(\mathbb{F}_{2},\mathbb{F}_{2}).$ The Kameko homomorphism is one of the useful tools to study the dimension of the domain of the Singer transfer. Singer conjectured that the algebraic transfer is always a monomorphism, but this remains open in general case. In this paper, by constructing a novel algorithm implemented in the computer algebra system OSCAR for computing $GL(q)$-invariants of the kernel of the Kameko homomorphism, we disprove Singer’s conjecture for bidegree $(6,6+36).$
@@
Adams spectral sequences; Steenrod algebra; Hit problem; Algebraic transfer; OSCAR (computer algebra system)
2020 MSC:
55Q45, 55S10, 55S05, 55T15, 68W30
@@ 1 Introduction and statement of the main outcome
Introduction. Let $\mathbb{F}_{2}$ be the prime field with two elements. We use the shorthand $H^{*}(X)$ (resp. $H_{*}(X)$) for the singular cohomology (resp. homology) groups with coefficients in $\mathbb{F}_{2}$. The Steenrod algebra $\mathscr{A}$ is the algebra of all stable cohomology operations over $\mathbb{F}_{2}$ and plays a fundamental role in Algebraic Topology, particularly in stable homotopy theory. A central problem in this field is computing the stable homotopy groups of spheres. Despite many profound results, this problem remains challenging and is far from being fully solved. Researchers have developed deep theories and practical tools to understand and compute these groups. One of the most useful tools is the Adams spectral sequence, which approximates the 2-primary stable homotopy groups of the sphere spectrum $\mathbb{S}^{0}$. Its input is the cohomology of the Steenrod algebra, ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})=\bigoplus_{r\geq 0}{\rm Ext}^{q,r}_{\mathscr{A}}(\widetilde{H}^{*}(\mathbb{S}^{0})=\mathbb{F}_{2},\mathbb{F}_{2}),$ where $q$ is the homological degree and $r$ is the internal degree. For a deeper understanding of $\mathrm{Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}),$ readers may refer to papers such as [20, 14, 4, 5, 15, 16]. Within the scope of this paper, another efficient instrument that we are especially interested in is the Singer algebraic transfer, proposed by Singer in 1989 [33]. Before delving into the details of the Singer transfer, we will recall some pertinent aspects.
@@
Let $\mathcal{V}^{q}$ denote a $q$-dimensional $\mathbb{F}_{2}$-vector space. Since $\mathbb{F}_{2}$ is a prime field of size two, $\mathcal{V}^{q}$ can be regarded as a rank-$q$ elementary abelian 2-group. It is well-known that $H^{*}(\mathcal{V}^{q})\cong S(\mathcal{V}_{*}^{q})$, the symmetric algebra of the dual space $\mathcal{V}_{*}^{q}\equiv H^{1}(\mathcal{V}^{q})$. We can choose $x_{1},x_{2},\ldots,x_{q}$ to be a basis of $H^{1}(\mathcal{V}^{q})$. In this case, $P_{q}:=H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}[x_{1},x_{2},\ldots,x_{q}]$, the connected $\mathbb{N}$-graded polynomial algebra on generators of degree $1,$ equipped with the canonical unstable algebra structure over $\mathscr{A}.$ By dualizing, the mod-2 homology $H_{*}(\mathcal{V}^{q})$ is a divided power algebra on $q$ generators. Let $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q})$ be the subspace of $H_{*}(\mathcal{V}^{q})$ consisting of all elements that are annihilated by all positive degree Steenrod operations. The group $GL(q)$ acts regularly on $\mathcal{V}^{q}$ and therefore on $P_{q}$ and $H_{*}(\mathcal{V}^{q}).$ This action commutes with that of the algebra $\mathscr{A}$ and so acts on $\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}$ and $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q}).$ With the idea that the structure of the Ext groups can be studied through modular invariant theory, Singer [33] formulated a homomorphism denoted as:
@@
The domain of $Tr_{q}(\mathbb{F}_{2})$ is closely related to the structure of the tensor product $\mathbb{F}_{2}\otimes_{\mathscr{A}}H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}.$ Indeed, we give $\mathbb{F}_{2}$ the trivial $\mathscr{A}$-module structure. That is, the unit in $\mathscr{A}$ acts as a unit, while $Sq^{k}(\mathbb{F}_{2})=0$ for any $k\geq 1.$ Let $\mathscr{A}^{>0}$ denote the positive degree part of $\mathscr{A},$ and put
@@
where $\mathscr{A}^{>0}\cdot P_{q}$ refers to the subspace of $P_{q}$ composed of all homogeneous polynomials of the form $\sum_{k\geq 1}Sq^{k}(f_{k}),$ with $Sq^{k}\in\mathscr{A}^{>0}$ and $f_{k}\in P_{q}.$ Note that
@@
To provide readers with the context related to the inequality above, we restate the following important fact, which has been mentioned in [32, Remark 2.3]: A key element in analyzing $(QP_{q})_{n}$ is the set of ”spike” monomials—monomials where every exponent is of the form $2^{m}-1$. A classical result establishes that spike monomials do not belong to $\mathscr{A}^{>0}\cdot P_{q}$ (see also [38]). Additionally, in [18], Mothebe constructed a rather involved recursive function to enumerate all spike monomials of degree $n=2^{q-1}-q$, namely
@@
Understanding the structure and computing the dimensions of $(QP_{q})_{n}$ and the invariant spaces $[(QP_{q})_{n}]^{GL(q)}$ are extremely difficult problems, if not impossible, even with modern computer algebra systems. The Peterson conjecture [21], which was proven by Wood [40], provides further insight into the graded vector space $QP_{q}$. This conjecture states that $QP_{q}$ is trivial in degrees $n$ if $\mu(n)>q$, where $\mu(n)$ denotes the minimal integer $\zeta$ for which $n$ can be written as $\sum_{1\leq j\leq\zeta}(2^{d_{j}}-1)$ for some positive integers $d_{j}$. In light of this result, we now focus on investigating the domain of the algebraic transfer when $\mu(n)\leq q$. Notably, the condition $\mu(n)\leq q$ is equivalent to the useful formulation $\alpha(n+q)\leq q$, where $\alpha(k)$ is the number of 1’s in the binary expansion of the integer $k$. This helps characterize the relevant “families” of $n$ that satisfy this condition.
Over the past nearly four decades, the Singer transfer and related aspects have been extensively studied by numerous authors (see, e.g., [1, 3, 7, 8, 9, 10, 11, 12, 13, 17, 19, 21, 22, 34, 35, 36, 24, 25, 26, 27, 28, 29, 30, 31, 32, 37, 38, 39, 40]). In particular, in [1], Boardman showed that $Tr_{3}(\mathbb{F}_{2})$ is also an isomorphism. Remarkably, Singer [33] proved that the algebraic transfer fails to be surjective in bidegree $(5,14),$ after which he proposed the following ensuing unsettled conjecture.
@@
To prove Theorem 1.4, we construct and implement a new algorithm in the OSCAR computer algebra system [41]. The algorithm computes an explicit basis for both the kernel of the Kameko homomorphism and the space $(QP_{q})_{n},$ as well as their corresponding invariants, for any $q$ and $n$ where $n-q$ is even. (Our previous algorithm in [31], implemented in SageMath, did not perform these basis and invariant computations for the kernel of the Kameko homomorphism.) Our reasoning for choosing OSCAR over SageMath for this implementation is detailed in Note 3.5(C) of Section 3. Furthermore, we used this new algorithm to verify previously known results, including those we computed by hand and those published by other authors (see, e.g., [3, 10, 27, 35, 37]). Our algorithm’s output is consistent with these established findings.
@@
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [36], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [22, 23]). As a limitation of this approach, the result in [36] remains unverified by modern computer algebra systems such as OSCAR, Magma, or SageMath. In fact, verifying all the manual computations (now considered outdated) in [36] on a computer algebra system is very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [30]). This implies that the manual computational methods in [36], while not novel, become impractical at higher ranks, which explains why Singer’s Conjecture 1.1 remained open until this work. As discussed above, we address this limitation for the $(q,n)=(6,36)$ case in Theorem 1.4, where we provide full computational verification via computer algebra systems.
@@
Taking a different approach to Conjecture 1.1, Nguyen Huu Viet Hung [10] proposed the concept of a critical element within ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$. Specifically, a non-zero element $\mathbf{u}$ in ${\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is called critical, if it satisfies two conditions: (i) $\mu(2n+q)=q$, and (ii) the image of $\mathbf{u}$ under the classical squaring operation $Sq^{0}$ is zero.
It is well-established that $Sq^{0}$ is a monomorphism in positive stems of ${\rm Ext}_{\mathscr{A}}^{q,q+n}(\mathbb{F}_{2},\mathbb{F}_{2})$ for $q<5,$ thereby implying the absence of any critical element for $q<5.$ Remarkably, Hưng’s work [10, Theorem 5.9] states that Singer’s Conjecture 1.1 is not valid, if the algebraic transfer detects critical elements.
@@
Following Hưng [12, Theorem 2.1], if a critical element $\widehat{\mathbf{u}}\in{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ is in the image of the transfer $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}),$ then $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ is not a monomorphism. By [12, Theorem 2.2], the existence of a positive stem critical element $\widehat{\mathbf{u}}\in{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ in the image of the transfer $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ is equivalent to the existence of a positivestem critical element $\mathbf{u}$ in the image of the transfer $Tr_{q+1}(\mathbb{F}_{2}).$ If the existences happen, then both $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ and $Tr_{q+1}(\mathbb{F}_{2})$ are not injective.
We know that the algebraic Kahn-Priddy homomorphism $t_{*}$ defined by
@@
is a surjection in positive stems (see [12]). In particular, the restriction of $t_{*}$ maps ${\rm Im}(Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}))$ onto ${\rm Im}(Tr_{q+1}(\mathbb{F}_{2}))$. Hence, for $q=5$ and $n=26$, there exists a non-zero element $\widehat{D_{2}}$ in the image of the transfer
$Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty}):(\mathbb{F}_{2}\otimes_{GL(5)}\mathcal{P}_{\mathscr{A}}(H_{*}\mathcal{V}^{5}\otimes\widetilde{H}_{*}\mathbb{R}P^{\infty}))_{26}\longrightarrow{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$
with $\widehat{D_{2}}\in{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ and $t_{*}(\widehat{D_{2}})=D_{2}\in{\rm Ext}^{6,6+26}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$. Consequently, by a completely analogous argument to that used for the elements $\widehat{Ph_{1}}$ and $\widehat{Ph_{2}}$ in [12], it may be concluded that $\widehat{D_{2}}$ is also a critical element and is not in the image of $Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$.
Theoretically, the approach to Singer’s Conjecture 1.1 via critical elements is promising. In practice, however, finding and characterizing these critical elements is computationally difficult.
@@
It is known, by Chen [6], that the following element $\widetilde{t}$ is a representative of $t$:
$\widetilde{t}$ $=\lambda_{5}\bigl(\lambda_{9}\lambda_{3}\lambda_{5}\lambda_{7}^{2}+\lambda_{6}\lambda_{0}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{3}\lambda_{5}\lambda_{1}\lambda_{15}\lambda_{7}\bigr)$
$+\lambda_{3}\bigl(\lambda_{8}\lambda_{0}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{6}\lambda_{2}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{5}\lambda_{9}\lambda_{5}\lambda_{7}^{2}+\lambda_{3}\lambda_{5}\lambda_{7}\lambda_{3}\lambda_{15}\bigr).$
Using this result together with Theorem 1.4 and our algorithm given in [31] for determining preimages in the lambda algebra, we hope that there will be an answer to Conjecture 1.7.